OpenStudy (anonymous):

Write the equation of the line which passes through (–4, 2) and is parallel to y = –x + 6 in slope-intercept form.

OpenStudy (anonymous):

y=-x -2

jimthompson5910 (jim_thompson5910):

y = -x+6 has a slope of -1 since it's in the form y = mx+b (and m is the slope) ANY parallel line to this line will have a slope of -1 also So the equation of this unknown parallel line is y = -x+b but since this line passes through (-4,2), we know that x = -4 and y = 2 So 2 = -(-4) + b 2 = 4 + b 2-4 = b -2 = b b = -2 So the equation of the parallel line that passes through (-4,2) is y = -x-2

OpenStudy (anonymous):

Write the equation of the line which passes through (2, –3) and is perpendicular to y = 4x + 7 in standard form.

jimthompson5910 (jim_thompson5910):

With perpendicular lines, the slopes are negative reciprocals of each other. So here the given slope is 4, which means that the perpendicular slope will be -1/4 (flip the fraction and the sign) So the perpendicular equation will look like \[\large y = -\frac{1}{4}x+b\] Use the fact that the line passes through (2,-3) and plug in x=2 and y = -3 into the equation above to solve for 'b' to get your answer

OpenStudy (anonymous):

Graciiiiiias